OpenStudy (anonymous):

Find the product of (3x + 7y)(3x − 7y).

OpenStudy (anonymous):

You going to want to use the distributive property for this one. or FOIL as it is usually called in classrooms.

OpenStudy (anonymous):

3x(3x)+(3x)(-7y)+(7y)(3x)+(7y)(-7y)

OpenStudy (anonymous):

(3x+7y)(3x-7y) by formula,(a+b)(a-b)=a*a-b*b (3x*3x)-(7y*7y)=9x to the power 2-49y to the power 2